This ones got me stumped
A circular cylinder has a volume of 108pie cubic feetand a radius of 3 feet. What is the height of the cylinder?

Answer by checkley71(8403)

108PI=PI*3^2*H
DIVIVING BY PI WE GET
108=9H
H=108/9
H=12 FOR THE HEIGHT.
PROOF
108PI=PI*9*12
108PI=108PI